# L2R: Low-Rank and Lipschitz-Controlled Routing for Mixture-of-Experts

### v1

@@ 4.2 Experiments with ViT on ImageNet
5 Conclusion
We introduced Low-rank & Lipschitz-controlled Routing (L2R), a routing framework that improves MoE specialization by jointly redesigning the routing space and scoring geometry. L2R replaces high-dimensional linear routing with a shared low-rank latent space, adopts Saturated Inner-Product Scoring (SIPS) to bound magnitude sensitivity while preserving controlled magnitude for adaptive expert weighting, and extends expressiveness with lightweight multi-anchor heads. Experiments on OLMoE pretraining and ViT-based ImageNet training show consistent gains, faster convergence, and balanced expert usage, indicating that L2R provides a practical and stable routing recipe across both language and vision MoE systems.
Impact Statement
Acknowledgment
This research was supported in part by AMED Grant Number JP256f0137006, and JSPS KAKENHI Grant Numbers JP23K11141, JP23K21676, JP24K02942, JP24K23849, and JP25K21218.
Appendix A Additional Geometry of Query-Magnitude Saturation
@@
Appendix B Gradient Sensitivity and Lipschitz Bounds of SIPS
@@
Appendix C Complexity and Efficiency
@@
Appendix D Optimization Objective Details
We train L2R under the standard optimization objective adopted by prior sparse MoE models, without introducing additional loss terms. Specifically, the objective consists of the task loss, a load-balancing loss to encourage uniform expert utilization, and the router z-loss (Zoph et al., 2022) used in OLMoE (Muennighoff et al., 2025) to stabilize gating logits in large-scale language model training.
@@ Appendix D Optimization Objective Details
Following OLMoE (Muennighoff et al., 2025), we additionally employ the router z-loss for large-scale language model training to regularize the magnitude of pre-softmax routing logits. Let $\bm{z}_{t}\in\mathbb{R}^{N}$ denote the routing logits for token $t$. The z-loss is defined as:
@@
Appendix E Experiments with OLMoE: Data Preparation
@@
Appendix F Training Configuration Details
F.1 Training Configuration for OLMoE Experiments
This appendix reports the core training hyperparameters for OLMoE experiments in Table 8. To isolate routing effects, we follow the original OLMoE (Muennighoff et al., 2025) settings and keep the backbone architecture, optimization recipe, and MoE auxiliary losses fixed across all router variants, modifying only router-specific components (e.g., projection rank, scoring, or head design) when applicable.
@@ F.1 Training Configuration for OLMoE Experiments
All models are trained from scratch on the 10B-token in-distribution dataset described in Appendix E. We follow the standard OLMoE recipe with a Transformer backbone of $d_{\text{model}}=2048$, $N=64$ experts per MoE block, and top-$k=8$ sparse activation. Optimization is performed using AdamW with cosine learning-rate decay, and all experiments are trained with Distributed Data Parallel (DDP) for scalability and stability. All runs are conducted on an $8\times$ NVIDIA B200 (Blackwell) setup.
The original OLMoE implementation relies on Megablocks (Gale et al., 2023) for MoE kernel acceleration. However, Megablocks currently does not support Blackwell sm_100 image configuration, which limits its compatibility on B200. To ensure correct execution and efficient expert dispatch, we replace the Megablocks MoE module with Tutel (Hwang et al., 2023), which offers architecture-agnostic kernels and stable performance under modern GPU backends. This change affects only the MoE dispatch backend and does not alter the routing logic or model formulation, ensuring fair and controlled comparisons across routing variants. For full reproducibility, we will release the complete configuration files and training scripts with the codebase.
F.2 Training Configuration for ViT Experiments
@@
Appendix G Full Experimental Results for OLMoE Experiments
@@
Appendix H Downstream Evaluation Protocol
We follow the downstream evaluation protocol used by OLMoE (Muennighoff et al., 2025) and OLMES (Gu et al., 2025) for reporting model quality. Specifically, we evaluate a fixed set of commonsense and knowledge benchmarks at checkpoints along training and report the average score (Overall) across all included tasks. All evaluations are performed using the same prompt format and normalization conventions as in the OLMoE/OLMES setup.
H.1 Tasks and Metrics
Table 11 summarizes the tasks and evaluation settings used in this work. We report task accuracy for all benchmarks. For MMLU (Hendrycks et al., 2021), we follow the standard multiple-choice formulation and report both the mean MMLU score and the four group scores (Humanities, Social Sciences, STEM, Other) in Appendix G.
Table 11: Summary of downstream evaluation. We adopt the OLMoE/OLMES evaluation conventions. CF = completion/close (generative) formulation; MCF = multiple-choice formulation; Norm indicates the answer normalization used by the evaluator. ARC-E refers to ARC-Easy.
@@ H.1 Tasks and Metrics
MMLU (Hendrycks et al., 2021) MCF (5-shot) char Accuracy
ARC-E (Clark et al., 2018) CF (0-shot) none Accuracy
BoolQ (Clark et al., 2019) CF (0-shot) none Accuracy
HellaSwag (Zellers et al., 2019) CF (0-shot) char Accuracy
PIQA (Bisk et al., 2020) CF (0-shot) char Accuracy
SciQ (Welbl et al., 2017) CF (0-shot) none Accuracy
H.2 Early-Stage Evaluation
Our language MoE models are trained from scratch up to 10B tokens. At this early pretraining stage, some downstream benchmarks commonly reported in OLMoE/OLMES (e.g., WinoGrande (Sakaguchi et al., 2020) and ARC-Challenge (Clark et al., 2018)) tend to exhibit high variance and unstable ranking across checkpoints, because the model has not yet formed sufficiently robust linguistic and commonsense representations. Including such tasks at 10B tokens can introduce noise that obscures routing-induced differences. Therefore, we focus on a compact set of tasks (Table 11) that are more stable and informative in the 10B-token regime, and we report an overall mean across these tasks for comparability.
H.3 Benchmark Descriptions
@@
Appendix I Visualizations for OLMoE
I.1 Visualizations for OLMoE Training Results
@@
I.2 Routing Geometry Visualizations in OLMoE
@@
I.3 Expert Usage Heatmaps

### v2

@@ Appendix C Ablation Study for ImageNet
Table 8: Ablation on multi-anchor heads $H$ with rank r=2r{=}2 on ImageNet-1K.
Head Top-1 (%) $\uparrow$ Top-5 (%) $\uparrow$
1 75.47 91.31
4 76.58 91.50
8 76.22 91.79
16 75.20 91.00
Ablation on Multi-Anchor Heads $H$.
Table 8 varies the head count with fixed r=2r{=}2. We observe the strongest performance at a moderate head count around H=4H{=}4; larger $H$ does not yield consistent gains and can slightly degrade accuracy. This suggests that, for ImageNet-1K with relatively limited MoE depth applied to only the last 4 layers, excessive anchor multiplicity may over-fragment routing regions and introduce overfitting, whereas a small number of anchors provides sufficient multi-mode coverage.
Appendix D Additional Geometry of Query-Magnitude Saturation
@@
Appendix E Gradient Sensitivity and Lipschitz Bounds of SIPS
@@
Appendix F Complexity and Efficiency
@@
[…]
Table 10: Inference overhead on OLMoE.
Model Prefill Latency (ms) $\downarrow$ Prefill Tok/s $\uparrow$ Decode Latency (ms/step) $\downarrow$ Decode Tok/s $\uparrow$
Linear 180.78 22657.39 43.79 22.84
L2R (Dot)-H1 181.94 22512.97 46.29 21.60
L2R (SIPS)-H1 187.08 21894.31 47.06 21.25
L2R (SIPS)-H16 182.36 22461.40 47.12 21.22
As shown in Table 10, L2R introduces only a small prefill overhead relative to Linear. For example, L2R (SIPS)-H16 is close to Linear in prefill latency (182.36 ms vs. 180.78 ms), while decode latency increases by about 3.3 ms/step. This is consistent with our complexity analysis: L2R is parameter-efficient, but theoretical savings do not automatically translate into end-to-end wall-clock gains without further kernel- and system-level optimization.
Appendix G Optimization Objective Details
We train L2R under the standard optimization objective adopted by prior sparse MoE models, without introducing additional loss terms. Specifically, the objective consists of the task loss, a load-balancing loss to encourage uniform expert utilization, and the router z-loss [35] used in OLMoE [19] to stabilize gating logits in large-scale language model training.
@@ Appendix G Optimization Objective Details
Following OLMoE [19], we additionally employ the router z-loss for large-scale language model training to regularize the magnitude of pre-softmax routing logits. Let $\bm{z}_{t}\in\mathbb{R}^{N}$ denote the routing logits for token $t$. The z-loss is defined as:
@@
Appendix H Experiments with OLMoE: Data Preparation
@@
Appendix I Training Configuration Details
I.1 Training Configuration for OLMoE Experiments
This appendix reports the core training hyperparameters for OLMoE experiments in Table 12. To isolate routing effects, we follow the original OLMoE [19] settings and keep the backbone architecture, optimization recipe, and MoE auxiliary losses fixed across all router variants, modifying only router-specific components (e.g., projection rank, scoring, or head design) when applicable.
@@ I.1 Training Configuration for OLMoE Experiments
All models are trained from scratch on the 10B-token in-distribution dataset described in Appendix H. We follow the standard OLMoE recipe with a Transformer backbone of $d_{\text{model}}=2048$, $N=64$ experts per MoE block, and top-$k=8$ sparse activation. Optimization is performed using AdamW with cosine learning-rate decay, and all experiments are trained with Distributed Data Parallel (DDP) for scalability and stability. All runs are conducted on an $8\times$ NVIDIA B200 (Blackwell) setup.
The original OLMoE implementation relies on Megablocks [12] for MoE kernel acceleration. However, Megablocks currently does not support Blackwell sm_100 image configuration, which limits its compatibility on B200. To ensure correct execution and efficient expert dispatch, we replace the Megablocks MoE module with Tutel [15], which offers architecture-agnostic kernels and stable performance under modern GPU backends. This change affects only the MoE dispatch backend and does not alter the routing logic or model formulation, ensuring fair and controlled comparisons across routing variants. For full reproducibility, we will release the complete configuration files and training scripts with the codebase.
I.2 Training Configuration for ViT Experiments
@@
Appendix J Full Experimental Results for OLMoE Experiments
@@
Appendix K Downstream Evaluation Protocol
We follow the downstream evaluation protocol used by OLMoE [19] and OLMES [13] for reporting model quality. Specifically, we evaluate a fixed set of commonsense and knowledge benchmarks at checkpoints along training and report the average score (Overall) across all included tasks. All evaluations are performed using the same prompt format and normalization conventions as in the OLMoE/OLMES setup.
K.1 Tasks and Metrics
Table 15 summarizes the tasks and evaluation settings used in this work. We report task accuracy for all benchmarks. For MMLU [14], we follow the standard multiple-choice formulation and report both the mean MMLU score and the four group scores (Humanities, Social Sciences, STEM, Other) in Appendix J.
Table 15: Summary of downstream evaluation. We adopt the OLMoE/OLMES evaluation conventions. CF = completion/cloze (generative) formulation; MCF = multiple-choice formulation; Norm indicates the answer normalization used by the evaluator. ARC-E refers to ARC-Easy.
@@
MMLU [14] MCF (5-shot) char Accuracy
ARC-E [7] CF (0-shot) none Accuracy
BoolQ [6] CF (0-shot) none Accuracy
HellaSwag [33] CF (0-shot) char Accuracy
PIQA [1] CF (0-shot) char Accuracy
SciQ [28] CF (0-shot) none Accuracy
K.2 Early-Stage Evaluation
Our language MoE models are trained from scratch up to 10B tokens. At this early pretraining stage, some downstream benchmarks commonly reported in OLMoE/OLMES (e.g., WinoGrande [25] and ARC-Challenge [7]) tend to exhibit high variance and unstable ranking across checkpoints, because the model has not yet formed sufficiently robust linguistic and commonsense representations. Including such tasks at 10B tokens can introduce noise that obscures routing-induced differences. Therefore, we focus on a compact set of tasks (Table 15) that are more stable and informative in the 10B-token regime, and we report an overall mean across these tasks for comparability.
K.3 Benchmark Descriptions
@@
Appendix L Limitation and Future Work
A limitation of our study is that the language experiments focus on an early-stage OLMoE-based pretraining regime. This setting is chosen to enable controlled router comparisons under identical data mixture, architecture, optimization schedule, and auxiliary routing losses. While the results consistently show improved routing geometry and downstream performance at 10B tokens, further scaling to longer pretraining budgets and larger MoE configurations remains an important direction for future work.
In addition, this work primarily evaluates L2R under representative language and vision MoE settings. Future work may further investigate how the optimal routing rank and number of anchors vary with model size, expert count, and training duration. Extending L2R to additional architectures, downstream adaptation scenarios, and more diverse deployment settings would also help better understand its generality and practical impact.
Appendix M Impact Statement
Appendix N Visualizations for OLMoE
N.1 Visualizations for OLMoE Training Results
@@
N.2 Routing Geometry Visualizations in OLMoE
@@
N.3 Expert Usage Heatmaps
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction describe L2R as a low-rank and Lipschitz-controlled routing framework, and the main claims are supported by the reported language and vision MoE experiments.
  5. Guidelines:
    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations and future work in Appendix L, including the early-stage OLMoE-based pretraining regime and the need to evaluate longer training, larger MoE configurations, and broader architectures.
  10. Guidelines:
    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper defines the SIPS scoring function in the main text and provides the corresponding gradient-sensitivity and Lipschitz-bound analysis in Appendix E.
  15. Guidelines:
    • The answer [N/A] means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper reports the model settings, routing variants, datasets, objectives, evaluation protocol, and hyperparameters in the main text and appendices.
  20. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide anonymized code as supplementary material for review, including implementation details and instructions to reproduce the main experiments. A de-anonymized version will be released upon acceptance.
  25. Guidelines:
    • The answer [N/A] means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Training configurations, routing hyperparameters, datasets, evaluation protocols, and implementation details are provided in the main text and Appendix I.
  30. Guidelines:
    • The answer [Yes] means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars because full from-scratch MoE training is computationally expensive. Instead, we report consistent trends across multiple routing variants, ablations, and both language and vision settings.
  35. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  39. Answer: [Yes]
  40. Justification: The paper reports the GPU types used for the language and vision experiments, and Appendix F provides efficiency analysis and timing measurements. We do not report exhaustive memory traces for every run, but the provided compute and timing information is sufficient to understand the resource requirements of the main experiments.
  41. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  42. 9.
    Code of ethics
  43. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  44. Answer: [Yes]
  45. Justification: The work studies MoE routing mechanisms using existing research datasets and does not involve human subjects, private data, or deceptive data collection.
  46. Guidelines:
    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  47. 10.
    Broader impacts
  48. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  49. Answer: [Yes]
  50. Justification: The paper discusses broader impacts in Appendix M. In this study, we propose a new routing framework for Mixture-of-Experts models, which may yield positive impacts, including improving conditional computation efficiency, reducing computational resource requirements, and potentially decreasing energy consumption associated with large-scale training and inference.
  51. Guidelines:
    • The answer [N/A] means that there is no societal impact of the work performed.
    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  52. 11.
    Safeguards
  53. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  54. Answer: [N/A]
  55. Justification: The paper does not release a new pretrained model, image generator, or scraped dataset with high misuse risk.
  56. Guidelines:
    • The answer [N/A] means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  57. 12.
    Licenses for existing assets
  58. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  59. Answer: [No]
  60. Justification: The paper cites the original sources of the models, datasets, and benchmarks used in the experiments, but does not explicitly list the licenses and terms of use for each asset.
  61. Guidelines:
    • The answer [N/A] means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  62. 13.
    New assets
  63. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  64. Answer: [N/A]
  65. Justification: The paper does not introduce a new dataset or pretrained model asset. The planned release is code for reproducing the proposed routing method.
  66. Guidelines:
    • The answer [N/A] means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  67. 14.
    Crowdsourcing and research with human subjects
  68. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  69. Answer: [N/A]
  70. Justification: The paper does not involve crowdsourcing or research with human subjects.
  71. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  72. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  73. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  74. Answer: [N/A]
  75. Justification: The paper does not involve human subjects, so IRB approval is not applicable.
  76. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  77. 16.
    Declaration of LLM usage
  78. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
  79. Answer: [N/A]
  80. Justification: The core method does not use LLMs as an important, original, or non-standard component. MoE language models are used only as experimental architectures for evaluating the proposed routing method.
  81. Guidelines:
    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.